OpenStudy (anonymous):

Solve the equation by graphing: x2 – 5x – 24 = 0. A. {4, 6} B. {-2, 12} C. {-4, 6} D. {-3, 8}

OpenStudy (anonymous):

Solve for x first

Directrix (directrix):

D. {-3, 8} x^2 – 5x – 24 = 0 (x-8)*(x+3)=0 x = 8 or x =- 3